# On the relationship between the Wasserstein distance and differences in life expectancy at birth

### v4

@@ 1 Introduction
To the best of my knowledge, the first scholars who applied the optimal transport theory framework to the field of demography are Oeppen et al. (2021) and Cilek et al. (2023). While Oeppen et al. (2021) discussed optimisation in life-saving models of mortality, Cilek et al. (2023) used the Wasserstein distance to investigate cause-specific mortality differences between border regions.
@@ 4.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard index - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 5(b) and 6.
@@
Figure 6: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the Non-overlap index (Jaccard index)
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 8(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 8(a)).
@@ 4.4 Cohort life tables and over time comparisons
This is not surprising because $W_{1}$ reflects the absolute differences between the $l(x)$ functions. In other words, the $e_{0}$ difference for a pair of countries cannot be larger than its $W_{1}$ Wasserstein distance. After taking into account comparisons over time, the correlation between the $W_{1}$ Wasserstein distance and the two alternative measures (KL-divergence and the Non-overlap index) remains strong. This is depicted in figure 13 and 14.
Figure 13: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the Non-overlap index (Jaccard index) using cohort data, women
Figure 14: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the Non-overlap index (Jaccard index) using cohort data, men
@@ 5 Conclusion
Finally, it is worth noting that other measures, such as the non-overlap index or the Kullback–Leibler divergence, can also be used to assess distributional differences. Our analysis shows that these measures are highly correlated with the $W_{1}$ Wasserstein distance and yield broadly similar conclusions. However, the Wasserstein distance provides a particularly elegant and intuitive way to interpret differences in age-at-death distributions, linking them directly to the framework of optimal transport.
@@ 6 Acknowledgements
A slightly different version of this paper was submitted to Demographic Research Formal Relationship section. The paper was rejected as its added value was deemed very limited but there was consensus about the technical correctness. Maybe someone else finds it interesting or useful. I would like to thank Laura Ann Cilek for inspiring conversations about the Wasserstein distance and for technical discussions on its implementation for comparing life table age-at-death distributions. I am also grateful to Pavel Griogoriev, Sebastian Klüsener, Jiaxin Shi, and Wen Su for helpful comments and suggestions.

### v5

@@ 1 Introduction
To the best of my knowledge, the first scholars who applied the optimal transport theory framework to the field of demography are Oeppen et al. (2021), Cilek et al. (2023), and Shang and Haberman (2025). While Oeppen et al. (2021) discussed optimisation in life-saving models of mortality, Cilek et al. (2023) used the Wasserstein distance to investigate cause-specific mortality differences between border regions and Shang and Haberman (2025) introduced an approach for forcasting age-at death distributions on the basis of CDFs which is related to the $W_{1}$ Wasserstein distance through equation 2 and 3.
@@ 3.4 Using the Wasserstein distance for cause-specific mortality comparisons
3.5 Comparing the $W_{1}$ Wasserstein distance to the non-overlap index
Several approaches exist for quantifying differences between age-at-death distributions. For instance, Shi et al. (2022) introduced the non-overlap index (also referred to as the Jaccard similarity index) to measure differences between age-at-death distributions. The measure is formally defined as,
$J(p,q)=\frac{\int_{-\infty}^{\infty}\min(p(x),q(x))\,dx}{\int_{-\infty}^{\infty}\max(p(x),q(x))\,dx},$ (24)
where $p(x)$ and $q(x)$ are two probability density functions. The index equals one when the two distributions are identical and zero when their supports are entirely disjoint, so that larger values indicate greater similarity. The corresponding Jaccard distance is JD=1J(p,q)JD=1-J(p,q).
As its name suggests, the non-overlap index quantifies distributional differences in terms of the shared mass between two probability densities. Since the $W_{1}$ Wasserstein distance reflects the absolute difference between the corresponding cumulative distribution functions both measure might respond similarly to a shift of probability mass toward older ages and thus, track each other closely in demographic applications.
To examine this relationship formally, it is convenient to express $J$ through the Total Variation (TV) distance. As shown by Moulton and Jiang (2018), $J$ can be defined as,
$J=\frac{1-\mathrm{TV}}{1+\mathrm{TV}},$ (25)
with
$\mathrm{TV}(p,q)=\frac{1}{2}\int_{-\infty}^{\infty}\lvert p(x)-q(x)\rvert\,dx.$ (26)
The equation 25 shows that $J$ and $\mathrm{TV}$ carry identical information. To connect $\mathrm{TV}$ to $W_{1}$, we consider the demographically relevant case in which $q(x)=p(x-\delta)$ is a pure location shift of $p$ by $\delta>0$ years along the age axis, while the shape of the distribution remains unchanged. The TV distance becomes then,
$\mathrm{TV}(\delta)=\frac{1}{2}\int_{-\infty}^{\infty}\lvert p(x)-p(x-\delta)\rvert\,dx.$ (27)
Applying a first-order Taylor expansion, $p(x-\delta)\approx p(x)-\delta\,p^{\prime}(x)$, gives,
$\lvert p(x)-p(x-\delta)\rvert\approx\delta\lvert p^{\prime}(x)\rvert$, (28)
so that
$\mathrm{TV}(\delta)\approx\frac{\delta}{2}\int_{-\infty}^{\infty}\lvert p^{\prime}(x)\rvert\,dx.$ (29)
For any unimodal density, $p^{\prime}(x)$ changes sign exactly once, at the modal age $M$. Hence, we have a rising part, where $p^{\prime}(x)>0$,
$\int_{-\infty}^{M}p^{\prime}(x)\,dx=p(M)-p(-\infty)=p(M)-0=p(M),$ (30)
and a falling part, where $p^{\prime}(x)<0$,
$\int_{M}^{\infty}-p^{\prime}(x)\,dx=-(p(\infty)-p(M))=-(0-p(M))=p(M).$ (31)
The total variation of $p^{\prime}(x)$ therefore equals twice the modal density: $\int_{-\infty}^{\infty}\lvert p^{\prime}(x)\rvert\,dx=2\,p(M)$. Substituting into (29) yields
$\mathrm{TV}(\delta)\approx\delta\cdot p(M).$ (32)
Under a pure location shift, it follows directly that,
$W_{1}(p,q)=\delta.$ (33)
Combining (32) and (33) with the exact identity (25) gives the following approximation for the Jaccard index in terms of $W_{1}$,
$J\approx\frac{1-W_{1}\cdot p(M)}{1+W_{1}\cdot p(M)}.$ (34)
Equation (34) implies that $J$ is a function of $W_{1}$ alone, with the modal density $p(M)$ acting as a scaling factor that depends on the shape of the $d(x)$ distribution.
Figure 5: Comparing the exact Jaccard index with the approximated Jaccard index
The plot compares the exact Jaccard index with the approximated version on the basis of Human Mortality Database life tables for women and men in the period 199020241990-2024. The analysis includes 1563 pairs (combinations of year, sex, and country). The correlation is not linear. Even if $p(M)$ were identical across all pairs, the relationship between $J_{\text{approx}}$ and $W_{1}$ would not be linear. It is a strictly convex decreasing curve that flattens as $W_{1}$ grows because the denominator $(1+W_{1}\cdot p(M))$ increases with $W_{1}$, compressing Jaccard values toward zero for large shifts.
This chain of relationships, $W_{1}=\delta$, $\mathrm{TV}\approx\delta\cdot p(M)$ from the Taylor expansion, and $J=(1-\mathrm{TV})/(1+\mathrm{TV})$ exactly, suggests that both measures closely track each other in demographic applications. The correlation is high whenever mortality improvements are predominantly location shifts of the $d(x)$ distribution, i.e., deaths are shifted toward older ages with relatively modest changes in the shape of the distribution. When shape changes such as the compression of mortality (rectangularization) are also present, the modal density $p(M)$ varies across comparisons, introducing an bias to the otherwise tight $W_{1}$ to $J$ relationship.
Additionally, distributional differences are often measured by the Kullback-Leibler (KL) divergence, defined as
$\mathrm{KL}(p,q)=\int_{-\infty}^{\infty}p(x)\log\frac{p(x)}{q(x)}\,dx.$ (35)
The KL divergence quantifies how much one probability distribution diverges from a reference distribution: it equals zero if and only if $p$ and $q$ are identical, and increases as the two distributions become more dissimilar. Unlike the non-overlap index, the KL divergence has no direct interpretation in terms of the geometry of the distributions. Rather than measuring how far apart the two distributions sit on the age axis, it measures the expected information gain when distinguishing $p$ from the reference $q$, and is therefore sensitive to differences in distributional shape.
@@ 4.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard distance - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 6(b) and 7.
@@
Figure 7: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the non-overlap index (Jaccard distance)
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 9(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 9(a)).
@@ 4.4 Cohort life tables and over time comparisons
This is not surprising because $W_{1}$ reflects the absolute differences between the $l(x)$ functions. In other words, the $e_{0}$ difference for a pair of countries cannot be larger than its $W_{1}$ Wasserstein distance. After taking into account comparisons over time, the correlation between the $W_{1}$ Wasserstein distance and the two alternative measures (KL-divergence and the non-overlap index) remains strong. This is depicted in figure 14 and 15.
Figure 14: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the non-overlap index (Jaccard distance) using cohort data, women
Figure 15: Comparing the $W_{1}$ Wasserstein distance with the Kullback-Leibler divergence and the non-overlap index (Jaccard distance) using cohort data, men
@@ 5 Conclusion
Finally, it is worth noting that other measures, such as the non-overlap index or the Kullback–Leibler divergence, can also be used to assess distributional differences. Our analysis shows that these measures are highly correlated with the $W_{1}$ Wasserstein distance and yield broadly similar conclusions. Why is this strong association so evident in the demographic setting of age-at-death distributions? Age-at-death distributions show a characteristic smooth and unimodal shape and tend to shift toward older ages as mortality declines. Although the shape is not strictly invariant, changes are typically dominated by location shifts rather than substantial alterations in dispersion or modality. Under the approximation that two such distributions differ primarily by a shift, the $W_{1}$ distance has a direct mathematical relationship with the non-overlap index. In this setting, both measures are effectively driven by the same underlying displacement parameter, providing a formal explanation for their high empirical correlation. Nonetheless, both measures should not be treated as interchangeable. The Wasserstein distance provides a particularly elegant and intuitive way to interpret differences in age-at-death distributions, linking them directly to the framework of optimal transport.
@@ 6 Acknowledgements
A slightly different version of this paper was submitted to Demographic Research Formal Relationship section. The paper was rejected as its added value was deemed very limited but there was consensus about the technical correctness. Maybe someone else finds it interesting or useful. I would like to thank Laura Ann Cilek for inspiring conversations about the Wasserstein distance and for technical discussions on its implementation for comparing life table age-at-death distributions. I am also grateful to Pavel Griogoriev, Sebastian Klüsener, Jiaxin Shi, and Wen Su for helpful comments and suggestions. I used AI for this work to assist with language editing, coding, and refining the presentation of mathematical arguments.